OpenStudy (anonymous):

Would I have to use the Double-Angle Identity for sin in this problem? Solve the equation 2sin(2x) = cos x

OpenStudy (campbell_st):

yes you have 4sinxcosx = cos x 4sinx = 1 sin x = 1/4 you may have to write it as a general solution as there doesn't appear to be a given domain for x

OpenStudy (anonymous):

\[4\cos(x)\sin(x)=\cos(x)\] \[4\cos(x)\sin(x)-\cos(x)=0\] \[cos(x)(4\sin(x)-1)=0\] \[\cos(x)=0, \sin(x)=\frac{1}{4}\]

OpenStudy (anonymous):

careful here. you cannot divide by cosine, because that gets rid of one of the solutions. one solution is where \(\cos(x)=0\)

OpenStudy (anonymous):

other you need \(\sin^{-1}(\frac{1}{4})\)

OpenStudy (anonymous):

Thank you